OpenStudy (sammie_akridge):

3x + y = 1 y + 4 = 5x What is the resulting equation?

OpenStudy (iluvsoccer):

So what are we doing? :/ Solving for something or what?

OpenStudy (blazeryder):

Substitution?

OpenStudy (sammie_akridge):

All it says is:For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 1 y + 4 = 5x What is the resulting equation?

OpenStudy (blazeryder):

Take a look at this: https://www.khanacademy.org/math/8th-engage-ny/engage-8th-module-4/8th-module-4-topic-d/v/the-substitution-method If you need more help just tag me!

OpenStudy (mbowers):

Arrange your x and y's on the same side first then you elimination by subtraction

OpenStudy (mbowers):

\[3x+y=1\] \[-5x +y =-4\] subtract \[3x+y=1\] \[5x-y=4\] now add 8x=5 divide by 8 [\x=\frac{ 5 }{ 8 }\]